# Generative Autoregressive Transformers for Model-Agnostic Federated MRI Reconstruction

### v1

@@ 5.1 Comparison Studies
In a first FL setup, we examined performance when individual sites employed entirely distinct types of model architectures (i.e., MoDL-3, rGAN, D5C5). Performance metrics are listed in Table 1 for single-coil and in Table 2 for multi-coil MRI datasets. For within-site reconstructions, FedGAT achieves significantly higher performance than all competing FL methods (p$<$0.05), except for FedGIMP that yields higher PSNR for R=8x on single-coil data at Site 1. On average, FedGAT offers (PSNR, SSIM) improvements of (2.06dB, 4.40%) over distillation baselines, and (2.07dB, 2.48)SSIM over generative baselines. Among FL methods, FedGAT attains the most competitive performance against the ‘Single’ benchmark, and outperforms the privacy-violating ‘Central’ by 0.42dB PSNR, 0.71% SSIM (p$<$0.05). For across-site reconstructions, FedGAT achieves significantly higher performance than all competing FL methods (p$<$0.05). On average, FedGAT offers (PSNR, SSIM) improvements of (1.22dB, 3.48%) over distillation baselines, and (2.60dB, 5.74)SSIM over generative baselines. In this setting, FedGAT attains the most competitive performance against the ‘Central’ benchmark while outperforming ‘Single’ by 1.59dB PSNR, 4.20% SSIM (p$<$0.05), which showcasing the generalization benefits of training models on multi-site datasets. Representative images reconstructed via FL methods in the examined model-heterogeneous setup are depicted in Fig. 3. Corroborating the quantitative assessments, FedGAT attains higher image quality compared to baselines, with sharper depiction of structural details, lower artifact and noise in reconstructed images. Meanwhile, competing FL methods suffer from a degree of spatial blur, residual noise and artifacts that lead to suboptimal depiction of detailed tissue structures.
@@
In a second FL setup, we examined performance when individual sites employed distinct variants of a common model type (i.e., MoDL with 3, 5, versus 7 cascades), resulting in varying levels of model complexity across sites. Naturally, the benefits of model-agnostic approaches can be lower in this setup, when compared against the first setup that prescribed entirely distinct model types. Still, we reasoned that FedGAT should continue to yield significant benefits over competing FL baselines. Performance metrics are listed in Table 3 for single-coil and in Table 4 for multi-coil MRI datasets. For within-site reconstructions, FedGAT achieves significantly higher performance than all competing FL methods, and attains the most competitive performances against the ‘Single’ and privacy-violating ‘Central’ (p$<$0.05), except for FedMD that yields higher PSNR on multi-coil data at Site 3. On average, FedGAT offers (PSNR, SSIM) improvements of (1.52dB, 1.15%) over distillation baselines, and (1.42dB, 3.32)SSIM over generative baselines. For across-site reconstructions, FedGAT achieves significantly higher performance than all competing FL methods (p$<$0.05). On average, FedGAT offers (PSNR, SSIM) improvements of (1.40dB, 0.99%) over distillation baselines, and (2.31dB, 4.29)SSIM over generative baselines. In this setting, FedGAT attains the most competitive performance against ‘Central’ while outperforming ‘Single’ by 1.42dB PSNR, 1.42% SSIM (p$<$0.05), indicating improved generalization. Representative images reconstructed in the second model-heterogeneous FL setup are depicted in Fig. 4. Corroborating the quantitative assessments, FedGAT attains higher image quality compared to baselines, with sharper depiction of structural details, lower artifacts and noise. Meanwhile, competing FL methods suffer from a degree of spatial blur, noise or artifacts that elicit degradations in depiction of detailed tissue structures.
@@
5.2 Ablation Studies
The primary aim of the current study is to enable collaborative training of generalizable reconstruction models with heterogeneous architectures across imaging sites. Towards this aim, FedGAT leverages synthetic MR images produced by a generative autoregressive transformer (GAT) as a privacy-preserving means of cross-site knowledge transfer. To examine the importance of the proposed GAT prior, we compared the fidelity of synthetic MR images generated by the GAT prior against those generated by adversarial (GAN) or diffusion (DDPM) priors. Fidelity of multi-site MR images generated by different priors as measured by FID are displayed in Fig. 5, and sample synthetic MR images are shown in Fig. 6. We find that GAT achieves the highest image fidelity, as manifested by its consistently lower FID scores compared to variants, and by its improved visual realism apparent in sample images. While GAN shows structural artifacts and suboptimal contrast and DDPM occasionally shows elevated noise levels, GAT achieves high image quality with lower artifacts and noise. We then assessed whether this improvement in synthetic image fidelity translates onto benefits in reconstruction performance. For this purpose, performance of reconstruction models trained on synthetic images from the GAT, GAN and DDPM priors were compared as listed in Table 5, in reference to a benchmark model that was trained using actual multi-site MR images. We find that the GAT prior yields significantly higher reconstruction performance than variants based on GAN and DDPM priors (p$<$0.05), and it attains the most competitive performance against the benchmark model. These results corroborate that the superior image fidelity of the proposed GAT prior translates onto improving the training of site-specific reconstruction models.
Next, we examined the efficacy of the knowledge-transfer approach in FedGAT that transfers information on the distribution of multi-site data via synthetic MR images. For this purpose, FedGAT was compared against a ‘w model agg.’ variant that simultaneously transferred knowledge and built reconstruction models via conventional model aggregation [36]. The comparisons were performed in model-homogeneous settings to permit use of model aggregation, and to avoid interactions between model architecture and data distribution across sites that could bias assessments. Performance metrics are listed in Table 6. FedGAT yields superior performance against ‘w model agg.’ in all cases ((p$<$0.05), except for within-site reconstructions at Site 3 where ‘w model agg.’ yields slightly higher SSIM. These results indicate that the decoupling of knowledge transfer and model building stages in FedGAT offers performance benefits against a coupled model aggregation procedure.
6 Discussion
Here we introduced a model-agnostic FL technique for MRI reconstruction that supports model-heterogeneous settings by decoupling cross-site knowledge transfer from building of reconstruction models. In particular, FedGAT mediates knowledge transfer among sites by sharing a global GAT prior that uses autoregressive predictions across multiple spatial scales to generate synthetic MR images, and a site prompt for precise control over the synthetic image distribution. Previous FL studies aiming to build generative priors have predominantly adopted adversarial approaches that can suffer from poor image quality due to training instabilities [20, 18, 74], or diffusion approaches that can suffer from extensive inference times and residual image noise [75, 76, 77, 78]. In comparison, the autoregressive nature of GAT offers improved image fidelity over adversarial and diffusion priors as indicated by our results. To curate a hybrid multi-site MRI dataset, the shared GAT prior is employed at each site to efficiently generate high-fidelity synthetic MRI datasets from remaining sites. Afterwards, site-specific reconstruction models are locally trained on this hybrid dataset to improve generalization while enjoying total freedom in architecture selection. Therefore, the proposed method holds great promise for advancing the practicality and scalability of FL in multi-institutional MRI studies.
@@ 6 Discussion
reconstruction PSNR $\uparrow$ SSIM $\uparrow$ PSNR $\uparrow$ SSIM $\uparrow$ PSNR $\uparrow$ SSIM $\uparrow$
@@
reconstruction PSNR $\uparrow$ SSIM $\uparrow$ PSNR $\uparrow$ SSIM $\uparrow$ PSNR $\uparrow$ SSIM $\uparrow$
@@
Table 6: Ablation study on the efficacy of the knowledge transfer approach in FedGAT.
@@
reconstruction PSNR $\uparrow$ SSIM $\uparrow$ PSNR $\uparrow$ SSIM $\uparrow$ PSNR $\uparrow$ SSIM $\uparrow$
FedGAT 4x 47.63$\pm$2.65 99.64$\pm$0.14 39.73$\pm$3.31 96.46$\pm$1.56 33.59$\pm$2.63 93.53$\pm$2.10
8x 41.60$\pm$2.85 98.78$\pm$0.54 34.58$\pm$3.36 94.14$\pm$2.33 28.85$\pm$2.62 87.75$\pm$3.34
w model agg. 4x 46.65$\pm$2.86 99.58$\pm$0.14 38.52$\pm$3.62 95.50$\pm$2.32 33.14$\pm$2.52 92.80$\pm$2.75
@@
reconstruction PSNR $\uparrow$ SSIM $\uparrow$ PSNR $\uparrow$ SSIM $\uparrow$ PSNR $\uparrow$ SSIM $\uparrow$
FedGAT 4x 36.71$\pm$3.01 95.02$\pm$1.96 40.28$\pm$2.65 96.39$\pm$1.81 43.52$\pm$2.83 98.06$\pm$1.04
8x 32.02$\pm$2.95 91.53$\pm$3.07 35.52$\pm$2.84 93.68$\pm$2.55 37.98$\pm$2.97 95.65$\pm$1.94
w model agg. 4x 35.83$\pm$3.12 94.15$\pm$2.54 39.89$\pm$2.70 96.19$\pm$1.95 42.59$\pm$3.26 97.54$\pm$1.64

### v2

@@ 5.1 Comparison Studies
In a first FL setup, we examined performance when individual sites employed entirely distinct types of model architectures (i.e., MoDL, rGAN, D5C5). Performance metrics are listed in Table 1 for coil-combined and in Table 2 for multi-coil MRI datasets. For within-site reconstructions, FedGAT achieves significantly higher performance than all competing FL methods (p$<$0.05), except for FedGIMP that yields higher PSNR for R=8x on coil-combined data at Site 1. On average, FedGAT offers (PSNR, SSIM) improvements of (2.1dB, 4.4%) over distillation baselines, and (2.1dB, 2.5%) over generative baselines. Among FL methods, FedGAT attains the most competitive performance against the ‘Single’ benchmark, and outperforms the privacy-violating ‘Central’ by 0.5dB PSNR, 0.7% SSIM (p$<$0.05). For across-site reconstructions, FedGAT achieves significantly higher performance than all competing FL methods (p$<$0.05). On average, FedGAT offers (PSNR, SSIM) improvements of (1.2dB, 3.5%) over distillation baselines, and (2.6dB, 5.7%) over generative baselines. In this setting, FedGAT attains the most competitive performance against the ‘Central’ benchmark while outperforming ‘Single’ by 1.6dB PSNR, 4.2% SSIM (p$<$0.05), which showcases the generalization benefits of training models on multi-site datasets. Representative images reconstructed via FL methods in the examined model-heterogeneous setup are depicted in Fig. 3. Corroborating the quantitative assessments, FedGAT attains higher image quality compared to baselines, with sharper depiction of structural details, lower artifact and noise levels in reconstructed images. Meanwhile, competing FL methods suffer from a degree of spatial blur, residual noise or artifacts that lead to suboptimal depiction of detailed tissue structures.
@@
In a second FL setup, we examined performance when individual sites employed distinct variants of a common model type (i.e., the unrolled MoDL model with 3, 5, versus 7 cascades), reflecting a scenario in which sites prefer varying levels of model complexity. Naturally, the benefits of model-agnostic approaches can be moderately lower in this setup, when compared against the first setup that prescribed entirely distinct model types. Still, we reasoned that FedGAT should continue to yield significant benefits over competing FL baselines. Performance metrics are listed in Table 3 for coil-combined and in Table 4 for multi-coil MRI datasets. For within-site reconstructions, FedGAT achieves significantly higher performance than all competing FL methods, and attains the most competitive performances against the ‘Single’ and privacy-violating ‘Central’ (p$<$0.05), except for FedMD that yields higher PSNR on multi-coil data at Site 3. On average, FedGAT offers (PSNR, SSIM) improvements of (1.6dB, 1.2%) over distillation baselines, and (1.5dB, 3.4%) over generative baselines. For across-site reconstructions, FedGAT achieves significantly higher performance than all competing FL methods (p$<$0.05). On average, FedGAT offers (PSNR, SSIM) improvements of (1.4dB, 1.0%) over distillation baselines, and (2.3dB, 4.3%) over generative baselines. In this setting, FedGAT attains the most competitive performance against ‘Central’ while outperforming ‘Single’ by 1.4dB PSNR, 1.4% SSIM (p$<$0.05), indicating improved generalization. Representative images reconstructed in the second model-heterogeneous FL setup are depicted in Fig. 4. Corroborating the quantitative assessments, FedGAT attains higher image quality compared to baselines, with sharper depiction of structural details, lower artifacts and noise. Meanwhile, competing FL methods suffer from a degree of spatial blur, noise or artifacts that elicit degradations in depiction of detailed tissue structures.
@@
Figure 5: Performance of site-specific reconstruction models (TransUNet) trained at a held-out site (Calgary) using a mixture of local and synthetic datasets. Synthetic datasets were generated by a GAT prior built in a three-site FL setup (BraTS, fastMRI, MIDAS). Separate models were trained while the percentage of samples from the local dataset was systematically varied. Results plotted for within-site reconstructions at the held-out site (Calgary), as well as across-site reconstructions in the FL sites (see legend).
@@
5.2 Generalization to a Non-Participating Site
Our comparison studies indicate that FedGAT enables the development of site-specific reconstruction models that generalize reliably across the participating sites in model-heterogeneous FL setups. Yet, FedGAT can also be deployed at a non-participating site that builds its reconstruction model by benefiting from the aggregated knowledge encapsulated by the GAT prior, enhancing within-site performance in scenarios where only limited local training data are available, and promoting generalization to datasets that diverge from the local data distribution. To assess potential benefits of FedGAT in a non-participating site, we considered a three-site FL setup to build the GAT prior (BraTS, fastMRI, MIDAS), and deployed it to train a transformer-based reconstruction model (TransUNet) at a held-out site (Calgary). To systematically examine the influence of local dataset size at the held-out site, separate models were trained using varying proportions of local samples and synthetic samples generated by the GAT prior. For reference, a 0% local / 100% synthetic setting (denoted 0%) indicates training solely on synthetic data, while a 100% local / 0% synthetic setting (denoted 100%) indicates training solely on local data; intermediate percentages reflect mixtures of the two sources.
Models trained at the held-out site were used to reconstruct images at the held-out site and also at each individual participating site, as shown in Fig. 5. For within-site reconstructions at the held-out site, while the model trained only on synthetic samples (0%) moderately underperforms compared to the reference model trained only on local samples (100%), models trained with mixed data perform competitively. On average, the performance gap between the 0% model and the reference 100% model is -0.8dB in PSNR and -6.2% in SSIM. In contrast, models trained with 40%–80% local data exhibit average differences of only +0.2dB in PSNR and +0.7% in SSIM relative to the 100% model. This result indicates that synthetic MR images produced by FedGAT can support training of performant reconstruction models at non-participating sites even when local training data are scarce. For across-site reconstructions at participating sites, the model trained only on local samples (100%) exhibits poor generalization performance relative to the model trained only on synthetic samples (0%). However, models trained with a mixture of local and synthetic data (i.e., 20%–80% local) show improved generalization, with an average gain of 6.5dB in PSNR and 19.6% in SSIM compared to the 100% model. This finding suggests that incorporating synthetic samples can help boost the generalization capability of models trained at non-participating sites to other imaging sites.
@@ 5.2 Generalization to a Non-Participating Site
[…]
@@
[…]
@@
Table 6: Susceptibility to information leakage for GAN, DDPM and GAT priors. FID scores of synthetic images were measured in the absence of (w/o int) and in the presence of (w int) an intruder site in the FL setup. $\Delta$(w/o int - w int) denotes the degradation level in image quality due to the intruder site.
Site 1 (BraTS) Site 2 (fastMRI) Site 3 (MIDAS) Average
w/o int w int w/o int w int w/o int w int $\Delta$(w/o int - w int)
GAN 7.47 14.05 1.15 2.77 2.68 6.43 3.98
DDPM 10.55 16.02 6.02 11.36 8.92 11.87 4.59
GAT 0.09 0.12 0.34 1.13 0.39 0.66 0.36
Table 7: Reconstruction performance of FedGAT variants built to examine the contribution of individual method components. Boldface indicates the top-performing variant in each task.
@@
[…]
FedGAT 4x 47.90$\pm$2.70 99.66$\pm$0.13 39.76$\pm$3.15 96.73$\pm$1.34 34.02$\pm$2.62 93.72$\pm$2.18
8x 41.61$\pm$2.88 98.66$\pm$0.60 35.34$\pm$3.05 94.60$\pm$2.12 29.56$\pm$2.53 89.35$\pm$3.29
w one-hot 4x 44.48$\pm$2.64 98.86$\pm$0.56 39.24$\pm$2.47 96.33$\pm$1.40 33.21$\pm$2.39 92.48$\pm$2.13
prompt 8x 38.00$\pm$2.64 87.56$\pm$5.25 34.35$\pm$2.58 92.30$\pm$1.87 28.96$\pm$2.47 85.56$\pm$3.07
w/o site 4x 46.70$\pm$2.95 99.53$\pm$0.18 38.99$\pm$3.79 95.95$\pm$2.03 33.45$\pm$2.67 92.83$\pm$2.73
prompt 8x 41.11$\pm$2.99 98.22$\pm$0.91 34.10$\pm$3.71 92.93$\pm$3.23 29.39$\pm$2.55 88.48$\pm$3.72
w/o scales 4x 45.66$\pm$2.89 99.43$\pm$0.20 39.12$\pm$3.20 96.49$\pm$1.42 33.30$\pm$2.66 93.13$\pm$2.11
8x 41.17$\pm$2.86 98.58$\pm$0.52 34.89$\pm$3.05 94.18$\pm$2.04 29.39$\pm$2.60 89.11$\pm$3.27
w/o autoreg 4x 42.95$\pm$2.44 96.25$\pm$2.22 38.79$\pm$2.23 96.01$\pm$1.26 31.51$\pm$2.42 87.11$\pm$3.68
8x 37.65$\pm$3.32 91.45$\pm$3.80 34.28$\pm$2.33 92.74$\pm$1.84 26.84$\pm$2.57 82.31$\pm$3.80
w model agg 4x 46.65$\pm$2.86 99.58$\pm$0.14 38.52$\pm$3.62 95.50$\pm$2.32 33.14$\pm$2.52 92.80$\pm$2.75
@@
[…]
FedGAT 4x 36.68$\pm$4.25 95.00$\pm$2.46 40.11$\pm$6.97 96.42$\pm$3.56 43.24$\pm$4.84 98.02$\pm$1.94
8x 32.09$\pm$4.00 91.60$\pm$3.85 35.28$\pm$6.81 93.82$\pm$5.38 38.25$\pm$4.40 96.43$\pm$2.79
w one-hot 4x 36.22$\pm$3.87 94.40$\pm$2.64 38.84$\pm$6.17 95.67$\pm$3.55 41.86$\pm$3.66 97.59$\pm$1.65
prompt 8x 31.66$\pm$3.69 88.93$\pm$4.22 33.48$\pm$5.19 86.56$\pm$4.41 36.18$\pm$3.19 89.93$\pm$4.59
w/o site 4x 36.34$\pm$4.33 94.50$\pm$2.75 39.99$\pm$6.85 96.18$\pm$3.86 42.85$\pm$4.85 97.67$\pm$2.33
prompt 8x 31.56$\pm$3.92 90.49$\pm$4.14 35.21$\pm$6.49 93.38$\pm$5.36 37.48$\pm$4.61 95.39$\pm$3.28
w/o scales 4x 34.16$\pm$3.88 91.25$\pm$3.97 38.14$\pm$6.92 94.76$\pm$5.29 42.11$\pm$4.55 97.50$\pm$2.28
8x 30.33$\pm$3.91 87.18$\pm$5.31 34.22$\pm$6.37 92.58$\pm$5.79 36.76$\pm$4.31 95.28$\pm$3.22
w/o autoreg 4x 35.15$\pm$4.32 91.56$\pm$5.23 37.23$\pm$6.21 91.68$\pm$5.49 40.87$\pm$3.13 96.13$\pm$1.81
8x 30.56$\pm$4.46 87.52$\pm$6.01 32.25$\pm$6.17 86.88$\pm$5.95 35.97$\pm$3.33 92.10$\pm$3.06
w model agg 4x 35.83$\pm$3.12 94.15$\pm$2.54 39.89$\pm$2.70 96.19$\pm$1.95 42.59$\pm$3.26 97.54$\pm$1.64
@@
w synthetic 4x 36.31$\pm$4.21 94.80$\pm$2.66 39.90$\pm$6.77 96.31$\pm$3.71 42.64$\pm$4.87 97.90$\pm$2.05
8x 31.55$\pm$3.82 91.28$\pm$4.04 34.76$\pm$6.18 93.53$\pm$5.64 37.32$\pm$4.40 95.55$\pm$3.23
5.3 Ablation Studies
5.3.1 The GAT Prior
FedGAT leverages synthetic MR images as a privacy-preserving means for cross-site knowledge transfer on local data distributions, and to enable collaborative training of generalizable reconstruction models in model-agnostic FL settings. To assess the utility of the proposed GAT prior in this context, we conducted a set of ablation studies comparing it against mainstream adversarial (GAN; [73]) and diffusion (DDPM; [84]) priors. Similar to the GAT prior, global GAN and DDPM priors were trained to capture the distribution of multi-site MRI data by conditioning them on a site prompt. First, we examined the fidelity of synthetic MR images produced by alternative priors in capturing the attributes of the underlying data distributions. Fidelity was measured via the FID metric between synthetic and actual MR images. FID values for GAT, GAN and DDPM priors are displayed in Fig. 6, and sample synthetic MR images are shown in Fig. 7. We find that GAT achieves the highest image fidelity, as manifested by its consistently lower FID scores compared to alternatives, and by its improved visual realism apparent in sample images. While GAN-generated images show some structural artifacts and suboptimal contrast and DDPM-generated images occasionally have residual noise, GAT achieves higher image quality with lower artifacts and noise.
Second, we assessed whether the improvement in synthetic image fidelity achieved by the GAT prior translates onto benefits in reconstruction performance. For this purpose, performance of reconstruction models trained solely on synthetic images from the GAT, GAN or DDPM priors were compared, in reference to a benchmark model that was trained using actual multi-site MR images. Comparisons were performed in model-homogeneous settings (with MoDL at all sites) to isolate the effect of synthetic data quality and avoid confounding interactions between model architecture and data distribution across sites. As listed in Table 5, reconstruction models trained using the GAT prior yield significantly higher performance than those trained using GAN and DDPM priors (p$<$0.05), and they attain the most competitive performance against the benchmark model. These results corroborate that the superior image fidelity of the proposed GAT prior helps train more performant reconstruction models.
As an FL technique, FedGAT communicates the parameters of the GAT prior instead of actual MRI data to mitigate potential risks in patient privacy. Yet, the prior might be susceptible to a level of information leakage across sites [24]. Accordingly, we comparatively evaluated GAT, GAN and DDPM priors in terms of their susceptibility to information leakage. For this purpose, an intruder site was included in the original three-site FL setup (BraTS, fastMRI, MIDAS), with local dataset for the intruder site taken as random white-Gaussian noise images. Adopting a recent method to assess information leakage in FL, we examined the influence of the intruder site on the fidelity of synthetic images generated at remaining sites as an indication of information leakage [85]. Table 6 lists fidelity of synthetic multi-site MR images generated based on GAT priors trained in the original three-site FL setup, and in the setup including the intruder site. We observe that GAN and DDPM priors experience a notable degree of loss in image fidelity, whereas the GAT prior shows relatively minute differences in fidelity. This result suggests that the GAT prior offers enhanced reliability against information leakage.
5.3.2 Other Key Design Elements
Next, we conducted another set of ablation studies to examine the contribution of key design elements in FedGAT to its performance. To assess the importance of utilizing a learnable site prompt via a gated MLP module, a ‘w one-hot prompt’ variant used an unlearned one-hot encoding of site index as the site prompt, and a ‘w/o site prompt’ variant entirely removed the site prompt from the input of the transformer module. To assess the importance of generating discrete token maps via a multi-scale autoregressive prediction, a ‘w/o scales’ variant ablated multi-scale generation of discrete token maps by setting $S=1$ (i.e., using a single spatial scale), and a ‘w/o autoreg’ variant ablated autoregressive prediction across scales by removing $f_{<S}$ from the input to the transformer module (i.e., predicting discrete token maps without knowledge of token maps at previous scales). To assess the importance of using hybrid local-synthetic datasets for cross-site knowledge transfer, a ‘w model agg’ variant was formed that simultaneously transferred knowledge and built reconstruction models via conventional model aggregation [37], and a ‘w synthetic’ variant was formed that built site-specific reconstruction models exclusively on synthetic datasets. The comparisons were again performed in model-homogeneous settings to avoid confounding interactions between model architecture and data distribution across sites. Performance metrics for variant models are listed in Table 7. FedGAT yields superior performance against the variant models in all cases (p$<$0.05). These results indicate that the individual design elements in FedGAT pertaining to the use of a gated-MLP driven site prompt, autoregressive and multi-scale generation of feature maps, and cross-site knowledge transfer via image synthesis all contribute significantly to its reconstruction performance.
6 Discussion
Here we introduced a model-agnostic FL technique for MRI reconstruction that supports model-heterogeneous settings by decoupling cross-site knowledge transfer from building of reconstruction models. In particular, FedGAT mediates knowledge transfer among sites by sharing a global GAT prior that uses autoregressive predictions across multiple spatial scales to generate synthetic MR images, and a site prompt for precise control over the synthetic image distribution. Previous imaging studies aiming to build generative priors have predominantly adopted adversarial approaches that can suffer from poor image quality due to training instabilities [19, 17, 86], or diffusion approaches that can suffer from extensive inference times and residual image noise [87, 88, 89, 90]. In comparison, the autoregressive nature of GAT can offer improved image fidelity over adversarial and diffusion priors in FL settings as indicated by our results. To curate a hybrid multi-site MRI dataset, the shared GAT prior is employed at each site to efficiently generate high-fidelity synthetic MRI datasets from remaining sites. Afterwards, site-specific reconstruction models are locally trained on this hybrid dataset to improve generalization while enjoying freedom in architecture selection. Therefore, the proposed method holds great promise for advancing the practicality and scalability of FL in multi-institutional MRI studies.
6.1 Site Specialization